# Evidence for Intrinsic Galaxy Alignments in Ellipticity Autocorrelations out to $100\,h^{-1}\,{\rm Mpc}$ from SDSS Galaxies with DESI Imaging

### v1

@@ 4.2 Comparison with Models
We perform the likelihood analysis to quantitatively evaluate the detection significance of II autocorrelations for the three SDSS galaxy samples. We compute the $\chi^{2}$ statistics for each measured statistic of $\xi_{+,\ell},\xi_{-,\ell},\widetilde{\xi}_{-,L}$. We use the $\ell=0,2$ multipoles for $\xi_{+,\ell}$ and $\xi_{-,\ell}$ and the $L=4$ multipole for $\widetilde{\xi}_{-,L}$, all of which were shown in Figures 1 and 2. The full covariance matrix is used to include the correlations between different bins. We have two fitting parameters, ${\boldsymbol{\theta}}=(b_{K}\sigma_{8},\sigma_{v})$. The analysis is performed over the scales adopted, $6\leq r\leq 100~{}[\,h^{-1}\,{\rm Mpc}]$. Namely, we have 6 data points for each of the multipoles. See Table 1 for the degree of freedom for each analysis. The data points used for the analysis are enclosed by the vertical lines in Figure 2.
Figure 3 shows the parameter constraints obtained from the LOWZ sample. The results for the LRG and CMASS samples are not shown here because the trend of the constraints and their interpretation are very similar to the LOWZ result. The green, orange, and blue contours are results with $\xi_{+,\ell},\xi_{-,\ell}$, and $\widetilde{\xi}_{-,L}$, respectively. We obtain similar constraints on $b_{K}$ from $\xi_{-,\ell}$ and $\widetilde{\xi}_{-,L}$. It is expected because they are the same statistics but expanded by different bases (see also Inoue et al., 2025). This demonstrates that the linear-level signal of the $\mathrm{II}(-)$ correlation is well compressed into the single $L=4$ multipole if it is expanded by the associated Legendre basis. Thus, this allows us to emphasize the first clear evidence of the ellipticity autocorrelation up to $\sim 100\,h^{-1}\,{\rm Mpc}$ via the appropriate choice of the basis and using the high-quality imaging data.
@@
The $\mathrm{II}(+)$ correlation is not sensitive to the FoG effect, and the best-fit $b_{K}$ value is lower than that from the $\mathrm{II}(-)$ correlation. Our model includes nonlinear RSD but still adopts the linear shape bias model, and overpredicts the amplitude of the $\mathrm{II}(+)$ correlation (Okumura et al., 2024). The nonlinear shape bias beyond the NLA model has been extensively studied (Vlah et al., 2020; Akitsu et al., 2021; Bakx et al., 2023; Matsubara, 2024; Chen & Kokron, 2024; Akitsu et al., 2023b). Accounting for such higher-order shape bias effects will enable us to predict the $\mathrm{II}(+)$ correlation and to use the measured IA correlation functions down to smaller scales, which will further enhance the science return from IA of galaxies.
@@
Finally, let us investigate the full signals of the ellipticity correlations, namely, jointly analyze the $\mathrm{II}(+)$ and $\mathrm{II}(-)$ correlation functions. Since the latter has a flexibility of choice for the basis, we consider the two data vectors, $(\xi_{+,\ell},\xi_{+,\ell^{\prime}})$ and $(\xi_{+,\ell},\widetilde{\xi}_{-,4})$ with ,=$\ell,\ell^{\prime}=\{0,2\}$ and construct the corresponding full covariance matrix. The constraints are shown in Table 1. As seen in Eqs. (17)–(20), all the ellipticity autocorrelation functions are derived from $P_{EE}$. Nevertheless, surprisingly, combining them leads to $\sim 10\%$ improvement for the detectability compared to the $\mathrm{II}(-)$-only constraints. Even in this case, the choice of the associated Legendre basis for the $\mathrm{II}(-)$ correlation provides almost the equivalent of the case of the standard Legendre basis that requires the measurement of the multipoles at least $\ell=0$ and 2.
@@
We have presented the first observational evidence for IA of galaxies in ellipticity autocorrelation functions extending to $100\,h^{-1}\,{\rm Mpc}$, using the LRG, LOWZ and CMASS galaxy samples from SDSS and SDSS-III BOSS over the redshift range $0.16\leq z\leq 0.70$. While such autocorrelations are significantly more difficult to measure than cross-correlations with galaxy density fields, they contain unique physical information, including potential imprints of gravitational waves, which are inaccessible to GI correlations.
To mitigate the projection effects of galaxy images along the line of sight, we expand the $\mathrm{II}(-)$ correlation in terms of associated Legendre polynomials, which isolate the linear-order information into a single multipole. This basis choice, combined with high-quality DESI imaging data, yields a clear power-law correlation extending to scales comparable to the BAO features. The results represent the first unambiguous evidence of large-scale shape autocorrelations across multiple redshift bins. Although the $\mathrm{II}(+)$ correlation is noisier, combining it with $\mathrm{II}(-)$ leads to a $\sim 10\%$ improvement in detection significance. This enhancement is notable, given that the two are mathematically equivalent under the linear alignment model.
Our analysis establishes a robust observational foundation for cosmological applications of ellipticity autocorrelations. These include future efforts to constrain the stochastic gravitational wave background as well as the expansion history of the universe. In particular, since BAO signatures are also encoded in IA statistics (Chisari & Dvorkin, 2013; Okumura et al., 2019; Xu et al., 2023b), combining IA measurements with galaxy clustering promises improved geometric constraints. The joint constraints on growth and expansion rates with galaxy IA will be presented in our upcoming paper (Okumura, 2025).
When we were finalizing the work based on the SDSS spectroscopy data and limited to the redshift of $0.16\leq z\leq 0.70$, the DESI DR1 data were publicly released (DESI Collaboration et al., 2025). It naturally extends this analysis over a wider redshift range. Additionally, optimizing the IA signal through luminosity weighting to brighter galaxies (e.g., Seljak et al., 2009) may further enhance sensitivity. These directions will be pursued in future work.
TO thanks Toshiki Kurita for the discussion on handling the DESI imaging data. TO acknowledges support from the Taiwan National Science and Technology Council under Grants Nos. NSTC 112-2112-M-001-034-, NSTC 113-2112-M-001-011- and NSTC 114-2112-M-001-004-, and the Academia Sinica Investigator Project Grant No. AS-IV-114-M03 for the period of 2025–2029. Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III web site is http://www.sdss3.org/.
@@ Appendix A Multipole Expansion of Power Spectra and Correlation Functions
In this Appendix, we provide some formulas used to compute the theoretical models of the ellipticity correlation functions but not presented in the body text. Full formulas including higher-order multipoles are provided in Okumura et al. (2024)
[…]
@@
where $L\geq m=4$. While the former with $\ell=0,2,4$ contain the linear-order contributions, only the lowest-order coefficient with $L=4$, we show the formulas only up to $\ell=2$ because the hexadecapole was not used in this work. In the latter, $\mathcal{\widetilde{Q}}_{EE,4}$, contains the contribution in linear theory. Using $p^{(n)}$, these coefficients are given by
@@
Substituting Eq. (A1) into Eqs. (19) and (20) gives our predictions for $\xi_{+,\ell}$ and $\widetilde{\xi}_{-,L}$.
@@
$\xi_{-,2}(r)=$ $-\frac{\sqrt{35}}{7}\widetilde{\xi}_{-,4}(r)-\frac{\sqrt{91}}{14}\widetilde{\xi}_{-,6}(r)+\frac{4\sqrt{1309}}{231}\widetilde{\xi}_{-,8}(r)-\frac{43\sqrt{5005}}{6006}\widetilde{\xi}_{-,10}(r)+\frac{5\sqrt{2002}}{546}\widetilde{\xi}_{-,12}(r)+\cdots\,,$ (A6)

### v2

@@ 4.2 Comparison with Models
We perform the likelihood analysis to quantitatively evaluate the detection significance of II autocorrelations for the three SDSS galaxy samples. We compute the $\chi^{2}$ statistics for each measured statistic of $\xi_{+,\ell},\xi_{-,\ell},\widetilde{\xi}_{-,L}$. We use the $\ell=0,2$ multipoles for $\xi_{+,\ell}$ and $\xi_{-,\ell}$ and the $L=4$ multipole for $\widetilde{\xi}_{-,L}$, all of which were shown in Figures 1 and 2.
We use the full covariance matrix to account for small correlations between different separation bins and between different statistics, although the II correlations are dominated by shot noise and thus primarily by the diagonal elements (see Fig. 3 of Chuang et al., 2022). We have two fitting parameters, ${\boldsymbol{\theta}}=(b_{K}\sigma_{8},\sigma_{v})$. The analysis is performed over the scales adopted, $6\leq r\leq 100~[\,h^{-1}\,{\rm Mpc}]$. Namely, we have 6 data points for each of the multipoles. See Table 1 for the degree of freedom for each analysis. The data points used for the analysis are enclosed by the vertical lines in Figure 2.
Figure 3 shows the parameter constraints obtained from the LOWZ sample. The results for the LRG and CMASS samples are not shown here because the trend of the constraints and their interpretation are very similar to the LOWZ result. The green, orange, and blue contours are results with $\xi_{+,\ell},\xi_{-,\ell}$, and $\widetilde{\xi}_{-,L}$, respectively. The best-fitting values of $b_{K}\sigma_{8}$ and the reduced $\chi^{2}$ are shown in Table 1. Note that, because we set $q=0$ in Equation (5), the definition of $b_{K}$ used here differs from that in the literature, and values are not directly comparable.
@@
As demonstrated in Okumura et al. (2024), the $\mathrm{II}(+)$ correlation is not sensitive to the FoG effect, and the best-fit $b_{K}$ value is lower than that from the $\mathrm{II}(-)$ correlation. Besides, $\xi_{+,\ell}$ are poorly fit by our model for the CMASS sample. By excluding the small-scale data of the quadrupole and adopting $r_{\min}=25\,h^{-1}\,{\rm Mpc}$, the reduced $\chi^{2}_{\min}$ is converged to $\chi^{2}_{\min}/\mathrm{dof}\sim 1$. Our model incorporates nonlinear RSD but retains the linear shape bias model, resulting in an overprediction of the amplitude of the $\mathrm{II}(+)$ correlation and a poor fit to the data on small scales. The nonlinear shape bias beyond the NLA model has been extensively studied (Vlah et al., 2020; Akitsu et al., 2021, 2023b; Bakx et al., 2023; Chen & Kokron, 2024; Matsubara, 2024). Accounting for such higher-order shape bias effects will enable us to predict the $\mathrm{II}(+)$ correlation better and to use the measured IA correlation functions down to smaller scales, which will further enhance the science return from IA of galaxies.
We obtain similar constraints on $b_{K}$ from $\xi_{-,\ell}$ and $\widetilde{\xi}_{-,L}$, which is expected because they are the same statistics but expanded by different bases (see also Inoue et al., 2025). This demonstrates that the linear-level signal of the $\mathrm{II}(-)$ correlation, when expanded by the associated Legendre basis, is well compressed into the single $L=4$ multipole. Thus, this allows us to emphasize the first clear evidence of the ellipticity autocorrelation up to $\sim 100\,h^{-1}\,{\rm Mpc}$ via the appropriate choice of the basis and using the high-quality imaging data. Note that our model of the $\mathrm{II}(-)$ correlation, when expanded in the associated Legendre basis, provides a good $\chi^{2}$ fit to all three samples.
@@
Finally, let us investigate the full signals of the ellipticity correlations, namely, jointly analyze the $\mathrm{II}(+)$ and $\mathrm{II}(-)$ correlation functions. Since the latter has a flexibility of choice for the basis, we consider the two data vectors $(\xi_{+,\ell},\xi_{+,\ell^{\prime}})$ and $(\xi_{+,\ell},\widetilde{\xi}_{-,4})$ with $\ell,\ell^{\prime}=\{0,2\}$ and construct the corresponding full covariance matrix. The constraints are shown in Table 1. As seen in Equations (17)–(20), all the ellipticity autocorrelation functions are derived from $P_{EE}$. Nevertheless, combining them leads to $\sim 10\%$ improvement in the detectability compared to the $\mathrm{II}(-)$-only constraints. Even in this case, the choice of the associated Legendre basis for the $\mathrm{II}(-)$ correlation provides almost the equivalent of the case of the standard Legendre basis that requires the measurement of the multipoles at least $\ell=0$ and 2.
@@
We have presented the first observational evidence for IA of galaxies in ellipticity autocorrelation functions extending to $100\,h^{-1}\,{\rm Mpc}$, using the LRG, LOWZ, and CMASS galaxy samples from SDSS and SDSS-III BOSS over the redshift range $0.16\leq z\leq 0.70$. While such autocorrelations are significantly more difficult to measure than the cross correlations with galaxy density fields, they contain unique physical information, including potential imprints of GWs, which are inaccessible to GI correlations.
To mitigate the projection effects of galaxy images along the line of sight, we expand the $\mathrm{II}(-)$ correlation in terms of associated Legendre polynomials, which isolate the linear-order information into a single multipole. This basis choice, combined with high-quality DESI imaging data, yields a clear power-law correlation extending to scales comparable to the BAO features. The results represent the first unambiguous evidence of large-scale shape autocorrelations across multiple redshift bins. Although the $\mathrm{II}(+)$ correlation is noisier, combining it with $\mathrm{II}(-)$ leads to a $\sim 10\%$ improvement in detection significance.
Our analysis establishes a robust observational foundation for cosmological applications of ellipticity autocorrelations. These include future efforts to constrain the stochastic GW background as well as the expansion history of the Universe. In particular, since BAO signatures are also encoded in IA statistics (Okumura et al., 2019; Xu et al., 2023b), combining IA measurements with galaxy clustering promises improved geometric constraints. Joint constraints on growth and expansion rates with galaxy IA will be presented in our upcoming Letter (Okumura, 2025).
When we were finalizing the work based on the SDSS spectroscopy data and limited to the redshift of $0.16\leq z\leq 0.70$, the DESI DR1 data were publicly released (DESI Collaboration et al., 2025). They naturally extend this analysis over a wider redshift range. Additionally, optimizing the IA signal through luminosity weighting to brighter galaxies (e.g., Seljak et al., 2009) may further enhance sensitivity. These directions will be pursued in future work.
T.O. thanks Toshiki Kurita for the discussion on handling the DESI imaging data. T.O. acknowledges support from the Taiwan National Science and Technology Council under grant Nos. NSTC 112-2112-M-001-034-, NSTC 113-2112-M-001-011- and NSTC 114-2112-M-001-004-, and the Academia Sinica Investigator Project grant No. AS-IV-114-M03 for the period of 2025–2029. Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the US Department of Energy Office of Science. The SDSS-III website is http://www.sdss3.org/.
@@ Appendix A Multipole Expansion of Power Spectra and Correlation Functions
In this Appendix, we provide some formulas used to compute the theoretical models of the ellipticity correlation functions but not presented in the body text. Full formulas, including higher-order multipoles, are presented in Okumura et al. (2024)
Adopting a simple Gaussian function to characterize the FoG damping, all the multipole moments of the power spectra are expressed in a factorized form with the angular dependence encoded by $p^{(n)}(\alpha)\equiv\int^{1}_{-1}d\mu_{\boldsymbol{k}}\mu_{\boldsymbol{k}}^{2n}e^{-\alpha\mu_{\boldsymbol{k}}^{2}}=\gamma(1/2+n,\alpha)\alpha^{-(1/2+n)}$. The multipole moments of the $EE$ power spectrum expanded in terms of the standard and associated Legendre polynomials are expressed as
@@
where $L\geq m=4$. While the former with $\ell=0,2,4$ contain the linear-order contributions, we show the formulas only up to $\ell=2$ because the hexadecapole was not used in this work. In the latter, only the lowest-order coefficient with $L=4$, $\mathcal{\widetilde{Q}}_{EE,4}$, contains the contribution in linear theory. Using $p^{(n)}$, these coefficients are given by
@@
Substituting Equations (A1) into Equations (19) and (20) gives our predictions for $\xi_{+,\ell}$ and $\widetilde{\xi}_{-,L}$.
@@
$\xi_{-,2}(r)=$ $-\frac{\sqrt{35}}{7}\widetilde{\xi}_{-,4}(r)-\frac{\sqrt{91}}{14}\widetilde{\xi}_{-,6}(r)+\frac{4\sqrt{1309}}{231}\widetilde{\xi}_{-,8}(r)-\frac{43\sqrt{5005}}{6006}\widetilde{\xi}_{-,10}(r)+\frac{5\sqrt{2002}}{546}\widetilde{\xi}_{-,12}(r)+\cdots\,.$ (A6)